# Shadows of generalised Hayward spacetimes : in vacuum and with plasma

### v1

@@ IV.3 Damour-Solodukhin Wormhole $00<\sigma<1,\kappa=$ 0 :
$\dddot{x}|_{x=x_{\rm ph}}$ $>0\quad\implies 0<\sigma<\frac{2}{3}\implies\text{photon sphere}$ (35)
[…]
@@
The second set of solutions to eq.(12), along with eq.(13), gives $x_{\rm ph}=3\sigma$ and $b^{2}=27\sigma^{2}$. This set of solutions are only valid when $\sigma\geq\frac{2}{3}$ since below this value, the photon sphere radius would be located inside the throat radius and thus will not be part of the manifold. The stability condition for this case reduces to
$\dddot{x}|_{x=x_{\rm ph}}=\frac{2(3\sigma-2)}{81\sigma^{5}},$ (37)
@@
$\dddot{x}|_{x=x_{\rm ph}}$ $>0\quad\implies\frac{2}{3}<\sigma<1\implies\text{photon sphere}.$ (38)
@@
The difference in the two branches of the solutions of eq.(12), as mentioned above, can be clearly seen from the effective potential plots shown in Fig.4(a), where the nature of the potential changes from a single to a double peak with the critical point being $\sigma=\frac{2}{3}$. The potential for the wormhole is plotted as a function of tortoise coordinate $x_{*}$ which is defined as $dx_{*}^{2}=\frac{dx^{2}}{f_{1}(x)f(x)}$ and ranges from (,)($\infty$,$\infty$)( - ∞ , ∞ ).
We now move on to study the effect of plasma on the shadow of this wormhole. For the homogeneous plasma profile, the variation of the shadow radius as a function of $\sigma$ and plasma parameter $k_{0}$ is shown in Fig.5(a). Large values of $x_{\rm sh}$ is seen for higher values of both $\sigma$ and $k_{0}$. To highlight the effect of plasma, we also plot the difference in shadow radius between the with and without plasma cases denoted by $\Delta x_{\rm sh}=x_{\rm sh}-x_{\rm sh}|_{k_{0}=0}$ as shown in Fig.5(b). For the case of non-homogeneous plasma, the $x_{\rm sh}$ and $\Delta x_{\rm sh}$ are shown in Fig.6(a) and Fig6(b), respectively. Unlike the homogeneous case, we observe that the shadow radius reduces as the plasma parameter $k_{x}$ increases.
[…]
@@
The metric functions are of the form $f_{1}(x)=1$ and $f(x)=1-\frac{2x^{2}}{x^{3}+2\kappa^{2}}$ for the Hayward wormhole spacetime. The only possible solution to eq.(12) and eq.(13), for this wormhole, is $f(x)=0$ and $b^{2}=x_{\rm ph}^{2}$. Hence for this case, the circular photon orbits are located at the throat, which is defined as the largest root to the cubic polynomial equation $f(x)=0$ i.e. $x^{3}+2\kappa^{2}-2x^{2}=0$. We plot the effective potential as a function of the tortoise coordinates shown in Fig7(a), which is characterized by a single peak for all values of $\kappa$. To understand the stability of photon orbits, we compute $\dddot{x}$ numerically and conclude that these orbits are unstable for the complete range of $\kappa$ as shown by the brown curve in Fig.(7(b)). The instability of the photon spheres is evident from the effective potentials as well since for the complete range of $\kappa$ have a single peak around the throat. Finally, we observe that the shadow radius reduces monotonically with increasing $\kappa$ as shown by the magenta curve in Fig7(b).
@@
For the case of homogeneous plasma, it can be checked from eq.(29) that the shadow radius is independent of the plasma parameter and is identical to the vacuum case. However, for the case of non-homogeneous plasma, the shadow radius decreases as the plasma parameter $k_{x}$ increases as shown in Fig.8(a). We also show the variation of $\Delta x_{\rm sh}=x_{\rm sh}-x_{\rm sh}|_{k_{x}=0}$ in Fig.8(b) with varying $\kappa$ and $k_{x}$ that highlights the deviation of shadow radius from the vacuum case in presence of non-homogeneous plasma.
Figure 9: (a) The plot shows the relative location of the two positive photon spheres with respect to the outer horizon. Only the larger of the two photon spheres lies outside the horizon and, hence, the only physically relevant solution for generating shadow. (b)The plot shows the shadow radius of the Hayward BH in blue along with the Schwarzschild BH shadow radius in violet for comparison. The brown curve represents $\dddot{x}$, which is positive for the complete range of parameters, indicating that the photon sphere is indeed unstable.
[…]
@@
The introduction of homogeneous plasma profile leads to increase in shadow radius with increasing $k_{0}$ as shown in Fig.(10(a)) while the reduction in the shadow radius is seen for the non-homogeneous plasma as the plasma parameter $k_{x}$ increases, shown in Fig.(11(a)). The difference in the shadow radius in the presence and absence of plasma is shown as a function of metric and plasma parameters in Fig.10(b) for the homogeneous plasma and in Fig.11(b) for the non-homogeneous case.
IV.6 Hayward-Damour-Solodukhin wormhole: $10<\sigma<$ 1, $0<\kappa\leq\frac{4}{3\sqrt{3}}$
This wormhole spacetime is the most general case spanning the entire range of ($\sigma,\kappa$). The metric functions for this class are given by $f_{1}(x)=1-\frac{2\sigma x^{2}}{x^{3}+2\sigma\kappa^{2}}$ and $f(x)=1-\frac{2x^{2}}{x^{3}+2\kappa^{2}}$. There are two branches of solutions for eq.(12) given by $f(x)=0$ and $-f_{1}^{\prime}(x)+\frac{2f_{1}(x)}{x}=0$. These solutions determine the radius of the circular photon orbits and the impact parameter is given by eq.(13) which reduces to $b^{2}=\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}$ for this case.
@@ IV.6 Hayward-Damour-Solodukhin wormhole: $10<\sigma<$ 1, $0<\kappa\leq\frac{4}{3\sqrt{3}}$
These equations have been solved numerically, and we observe that in this wormhole geometry, there are either one, two, or three photon spheres depending on the value of metric parameters. We have plotted the number of photon orbits as a function of parameters in Fig.12. Note that for certain choices of parameters there are no photon spheres denoted by the black region of Fig.12. In the case of single-peak potential, the photon sphere is located near the throat, and for the double-peak potential, the photon orbits at the throat is an anti-photon sphere. For the triple peak potential, there are two photon spheres out of which one is located at the throat, and the two photon spheres are separated by an anti-photon sphere. To get a better visual representation of these scenarios, we have plotted the effective potential for specific parameters to cover all the possible cases in Fig.13(a). We compute the shadow radius corresponding to the largest of the photon spheres in case there are more than one photon sphere. The shadow radius for the complete parameter range of $\sigma$ and $\kappa$ is plotted in Fig.13(b). Larger shadow radius is observed for higher $\sigma$ and for a particular choice of $\kappa$.
@@
Finally, we move on to study the effect of plasma on the shadow of this wormhole metric. We plot the shadow radius and their difference with and without plasma for both homogeneous and non-homogeneous profiles in Fig.14 and 15 respectively. Since the variation of multiple parameters and their effect on shadow cannot be studied simultaneously, we choose certain representative parameter sets and study the effect of plasma on their shadows. For the homogeneous plasma profile, we observe that the shadow radius increases as the plasma parameter $k_{0}$ increases as shown in Fig.14(a). However, for the case of non-homogeneous plasma, the shadow radius decreases as the plasma parameter $k_{x}$ reduces as shown in Fig.15(a). To complete the analysis we also plot the difference in the shadow radius with and without plasma for both homogeneous and non-homogeneous profiles as shown in Fig.14(b) and 15(b) respectively.
@@
Schwarzschild BH $(1,-)$ $(0,-)$ ($\uparrow$, $\uparrow$) (($\uparrow$( ↑$\downarrow$)
Schwarzschild WH $(0,-)$ $(0,-)$ $(k_{0},-)$ ($\uparrow$,$\downarrow$)
DS WH (,)($\uparrow$,$\uparrow$)( ↑ , ↑ ) $(0,-)$ ($\uparrow$,$\uparrow$ , ↑) ($\uparrow$,$\downarrow$ , ↓)
Hayward WH $(0,-)$ (,)($\uparrow$,$\downarrow$)( ↑ , ↓ ) $(k_{0},-)$ (,)($\uparrow$,$\downarrow$)( ↑ , ↓ )
Hayward BH $(1,-)$ (,)($\uparrow$,$\downarrow$)( ↑ , ↓ ) (,)($\uparrow$,$\uparrow$)( ↑ , ↑ ) (,)($\uparrow$,$\downarrow$)( ↑ , ↓ )
Case VI WH (,)($\uparrow$,$\uparrow$)( ↑ , ↑ ) (,)($\uparrow$,$\downarrow$)( ↑ , ↓ ) (,)($\uparrow$,$\uparrow$)( ↑ , ↑ ) (,)($\uparrow$,$\downarrow$)( ↑ , ↓ )
@@
Table 2: Table shows the summary of the results obtained in our study. For each spacetime, the dependence of the shadow radius ($x_{\rm sh}$) on the metric and plasma parameters is shown for vacuum and in presence of homogeneous and non-homogeneous plasma profiles. When studying the dependence of $x_{\rm sh}$ on a particular parameter, the other parameters are kept fixed. The (,)($\uparrow$,$\downarrow$)( ↑ , ↓ ) indicates increasing and decreasing values respectively. In some cases, the $x_{\rm sh}$ is constant and does not depend on the parameters and are denoted by ‘--’.
V Conclusion

### v2

@@ 4.3 Damour-Solodukhin Wormhole $0<\sigma<1,\kappa=0$ :
$\dddot{x}|_{x=x_{\rm ph}}$ $>0\quad\implies 0<\sigma<\frac{2}{3}\implies\text{photon sphere}$ (4.4)
$\dddot{x}|_{x=x_{\rm ph}}$ $<0\quad\implies\frac{2}{3}<\sigma<1\implies\text{anti-photon sphere}$ (4.5)
@@
The second set of solutions to eq.(3.11), along with eq.(3.12), gives $x_{\rm ph}=3\sigma$ and $b^{2}=27\sigma^{2}$. This set of solutions are only valid when $\sigma\geq\frac{2}{3}$ since below this value, the photon sphere radius would be located inside the throat radius and thus will not be part of the manifold. The stability condition for this case reduces to
$\dddot{x}|_{x=x_{\rm ph}}=\frac{2(3\sigma-2)}{81\sigma^{5}},$ (4.6)
@@
$\dddot{x}|_{x=x_{\rm ph}}$ $>0\quad\implies\frac{2}{3}<\sigma<1\implies\text{photon sphere}.$ (4.7)
@@
The difference in the two branches of the solutions of eq.(3.11), as mentioned above, can be clearly seen from the effective potential plots shown in Fig.(4(a)), where the nature of the potential changes from a single to a double peak with the critical point being $\sigma=\frac{2}{3}$. The potential for the wormhole is plotted as a function of tortoise coordinate $x_{*}$ which is defined as $dx_{*}^{2}=\frac{dx^{2}}{f_{1}(x)f(x)}$ and ranges from $(-\infty,\infty)$.
We now move on to study the effect of plasma on the shadow of this wormhole. For the homogeneous plasma profile, the variation of the shadow radius as a function of $\sigma$ and plasma parameter $k_{0}$ is shown in Fig.(5(a)). Large values of $x_{\rm sh}$ is seen for higher values of both $\sigma$ and $k_{0}$. To highlight the effect of plasma, we also plot the difference in shadow radius between the with and without plasma cases denoted by $\Delta x_{\rm sh}=x_{\rm sh}-x_{\rm sh}|_{k_{0}=0}$ as shown in Fig.(5(b)). For the case of non-homogeneous plasma, the $x_{\rm sh}$ and $\Delta x_{\rm sh}$ are shown in Fig.(6(a)) and Fig(6(b)), respectively. Unlike the homogeneous case, we observe that the shadow radius reduces as the plasma parameter $k_{x}$ increases.
4.4 Hayward Wormhole: $\sigma=0\;,\;0<\kappa\leq\frac{4}{3\sqrt{3}}$
@@
The metric functions are of the form $f_{1}(x)=1$ and $f(x)=1-\frac{2x^{2}}{x^{3}+2\kappa^{2}}$ for the Hayward wormhole spacetime. The only possible solution to eq.(3.11) and eq.(3.12), for this wormhole, is $f(x)=0$ and $b^{2}=x_{\rm ph}^{2}$. Hence for this case, the circular photon orbits are located at the throat, which is defined as the largest root to the cubic polynomial equation $f(x)=0$ i.e. $x^{3}+2\kappa^{2}-2x^{2}=0$. We plot the effective potential as a function of the tortoise coordinates shown in Fig.(7(a)), which is characterized by a single peak for all values of $\kappa$. To understand the stability of photon orbits, we compute $\dddot{x}$ numerically and conclude that these orbits are unstable for the complete range of $\kappa$ as shown by the brown curve in Fig.(7(b)). The instability of the photon spheres is evident from the effective potentials as well since for the complete range of $\kappa$ have a single peak around the throat. Finally, we observe that the shadow radius reduces monotonically with increasing $\kappa$ as shown by the magenta curve in Fig.(7(b)).
@@
For the case of homogeneous plasma, it can be checked from eq.(3.28) that the shadow radius is independent of the plasma parameter and is identical to the vacuum case. However, for the case of non-homogeneous plasma, the shadow radius decreases as the plasma parameter $k_{x}$ increases as shown in Fig.(8(a)). We also show the variation of $\Delta x_{\rm sh}=x_{\rm sh}-x_{\rm sh}|_{k_{x}=0}$ in Fig.(8(b)) with varying $\kappa$ and $k_{x}$ that highlights the deviation of shadow radius from the vacuum case in presence of non-homogeneous plasma.
Figure 9: (a) The plot shows the relative location of the two positive photon spheres with respect to the outer horizon. Only the larger of the two photon spheres lies outside the horizon and, hence, the only physically relevant solution for generating a shadow. (b)The plot shows the shadow radius of the Hayward BH in blue along with the Schwarzschild BH shadow radius in violet for comparison. The brown curve represents $\dddot{x}$, which is positive for the complete range of parameters, indicating that the photon sphere is indeed unstable.
4.5 Regular Hayward black hole: $\sigma=1,\;0<\kappa\leq\frac{4}{3\sqrt{3}}$
Setting $\sigma=1,\;0<\kappa\leq\frac{4}{3\sqrt{3}}$ reduces the metric in eq.(2.3) to that of the regular Hayward BH case which was the original metric to be deformed to give rise to the generalized metric. For this spacetime, all the curvature scalars remain finite for the complete range of the manifold and hence it is considered as a regular BH solution. The metric functions reduce to $f_{1}(x)=f(x)=1-\frac{2x^{2}}{x^{3}+2\kappa^{2}}$ and is characterized with two horizons corresponding to the roots of $g_{tt}=0$. Solving eq.(3.11) for the photon sphere radius results in a sixth-order polynomial equation, which has been solved numerically. The polynomial equation has only two positive roots ($x_{\rm ph1},x_{\rm ph2}$), of which, we are interested in the one outside the outer horizon. Fig.(9(a)) shows the position of ($x_{\rm ph1},x_{\rm ph2}$) with respect to the outer horizon. The larger of the two photon spheres, i.e. $x_{\rm ph2}$, denoted by the brown curve in Fig.(9(a)), lies outside the outer horizon and hence gives rise to the shadow. We then calculate the $\dddot{x}$ for the photon orbit corresponding to $x_{\rm ph2}$ shown in brown curve in Fig.(9(b)) and conclude that they are all unstable. The corresponding shadow radius is shown as the blue curve in Fig.(9(b)) where the Schwarzschild shadow radius (magenta curve) is also shown for comparison.
@@ 4.5 Regular Hayward black hole: $\sigma=1,\;0<\kappa\leq\frac{4}{3\sqrt{3}}$
The introduction of homogeneous plasma profile leads to increase in shadow radius with increasing $k_{0}$ as shown in Fig.(10(a)) while the reduction in the shadow radius is seen for the non-homogeneous plasma as the plasma parameter $k_{x}$ increases, shown in Fig.(11(a)). The difference in the shadow radius in the presence and absence of plasma is shown as a function of metric and plasma parameters in Fig.(10(b)) for the homogeneous plasma and in Fig.(11(b)) for the non-homogeneous case.
4.6 Hayward-Damour-Solodukhin wormhole: $0<\sigma<1$, $0<\kappa\leq\frac{4}{3\sqrt{3}}$
This wormhole spacetime is the most general case spanning the entire range of ($\sigma,\kappa$). The metric functions for this class are given by $f_{1}(x)=1-\frac{2\sigma x^{2}}{x^{3}+2\sigma\kappa^{2}}$ and $f(x)=1-\frac{2x^{2}}{x^{3}+2\kappa^{2}}$. There are two branches of solutions for eq.(3.11) given by $f(x)=0$ and $f_{1}^{\prime}(x)+\frac{2f_{1}(x)}{x}=0$. These solutions determine the radius of the circular photon orbits and the impact parameter is given by eq.(3.12) which reduces to $b^{2}=\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}$ for this case.
@@ 4.6 Hayward-Damour-Solodukhin wormhole: $0<\sigma<1$, $0<\kappa\leq\frac{4}{3\sqrt{3}}$
These equations have been solved numerically, and we observe that in this wormhole geometry, there are either one, two, or three photon spheres depending on the value of metric parameters. We have plotted the number of photon orbits as a function of parameters in Fig.(12). Note that for certain choices of parameters there are no photon spheres denoted by the black region of Fig.(12). In the case of single-peak potential, the photon sphere is located near the throat, and for the double-peak potential, the photon orbits at the throat is an anti-photon sphere. For the triple peak potential, there are two photon spheres out of which one is located at the throat, and the two photon spheres are separated by an anti-photon sphere. To get a better visual representation of these scenarios, we have plotted the effective potential for specific parameters to cover all the possible cases in Fig.(13(a)). We compute the shadow radius corresponding to the largest of the photon spheres in case there are more than one photon sphere. The shadow radius for the complete parameter range of $\sigma$ and $\kappa$ is plotted in Fig.(13(b)). Larger shadow radius is observed for higher $\sigma$ and for a particular choice of $\kappa$.
@@
Finally, we move on to study the effect of plasma on the shadow of this wormhole metric. We plot the shadow radius and their difference with and without plasma for both homogeneous and non-homogeneous profiles in Fig.(14) and Fig.(15) respectively. Since the variation of multiple parameters and their effect on shadow cannot be studied simultaneously, we choose certain representative parameter sets and study the effect of plasma on their shadows. For the homogeneous plasma profile, we observe that the shadow radius increases as the plasma parameter $k_{0}$ increases as shown in Fig.(14(a)). However, for the case of non-homogeneous plasma, the shadow radius decreases as the plasma parameter $k_{x}$ reduces as shown in Fig.(15(a)). To complete the analysis we also plot the difference in the shadow radius with and without plasma for both homogeneous and non-homogeneous profiles as shown in Fig.(14(b)) and Fig.(15(b)) respectively.
@@
Schwarzschild BH (1,)(1,-) (0,)(0,-) ($\uparrow$, $\uparrow$) $(\uparrow$$\downarrow$)
Schwarzschild WH (0,)(0,-) (0,)(0,-) $(k_{0},-)$ ($\uparrow$$\downarrow$)
DS WH $(\uparrow,\uparrow)$ (0,)(0,-) ($\uparrow,\uparrow$) ($\uparrow,\downarrow$)
Hayward WH (0,)(0,-) $(\uparrow,\downarrow)$ $(k_{0},-)$ $(\uparrow,\downarrow)$
Hayward BH (1,)(1,-) $(\uparrow,\downarrow)$ $(\uparrow,\uparrow)$ $(\uparrow,\downarrow)$
Case VI WH $(\uparrow,\uparrow)$ $(\uparrow,\downarrow)$ $(\uparrow,\uparrow)$ $(\uparrow,\downarrow)$
@@
Table 2: Table shows the summary of the results obtained in our study. For each spacetime, the dependence of the shadow radius ($x_{\rm sh}$) on the metric and plasma parameters is shown for vacuum and in presence of homogeneous and non-homogeneous plasma profiles. When studying the dependence of $x_{\rm sh}$ on a particular parameter, the other parameters are kept fixed. The $(\uparrow,\downarrow)$ indicates increasing and decreasing values respectively. In some cases, the $x_{\rm sh}$ is constant and does not depend on the parameters and are denoted by ‘-’.
5 Constraints from Sgr $\rm A^{*}$
In the previous section, we conducted a general analysis of the shadows cast by the new spacetime solutions as seen by an asymptotic observer. In this section, we contextualize these results using observational data from EHT to determine if any of these spacetimes are viable candidates to represent an astrophysical compact object. Specifically, we consider the most recent EHT estimate for the angular diameter ($\Phi$) of the shadow of Sgr $\rm A^{*}$, given by $\Phi=48.7\pm 7.0\ \mu as$ [3] and study if any of these spacetimes can represent this object. This analysis incorporates priors on the distance to Sgr $\rm A^{*}$ and estimated BH mass which are obtained from: Keck and the Very Large Telescope Interferometer (VLTI) [72, 73]. The Keck team reported a distance of $r_{o}=7959\pm 59\pm 32\;pc$ and a BH mass of $M=(3.975\pm 0.058\pm 0.026)\times 10^{6}M_{\odot}$ [74]. In contrast, VLTI, in collaboration with GRAVITY estimates the distance to be $r_{o}=8277\pm 9\pm 33\;pc$ and mass $M=(4.297\pm 0.012\pm 0.040)\times 10^{6}M_{\odot}$ [72, 73].
In order to relate these results to our framework, we need to convert the angular shadow diameter to shadow radius. This is achieved by first relating the angular diameter of the shadow with shadow radius which is given by
$\frac{r_{\text{sh}}}{r_{o}}=\tan\left(\frac{1}{2}\Phi\right).$ (5.1)
Note that the distance $r_{o}$ used above is obtained from the priors considered for generating the angular diameter of the shadow. Also, the shadow radius obtained above is not dimensionless, it can be converted to a dimensionless quantity by normalizing it with appropriate factors of BH mass (from the mass prior), speed of light and the gravitational constant. Carrying out this computations using the VLTI and Gravity collaboration data we obtain the following constraint on dimensionless shadow radius
$4.35548\leq x_{\text{sh}}\leq 5.81695.$ (5.2)
In principle, the same analysis can be done with the Keck data, which would lead to marginally different bounds. However, the overall conclusions remain largely unaffected.
Figure 16: (a) The figure shows the shadow radius of Damour-Solodukin WH as a function of the parameter $\sigma$. Values of $\sigma$ close to 1 lie within the EHT constraint whereas the Schwarzschild BH shadow lies well within the bound. (b) The figure shows the shadow radius of Hayward WH as a function of the parameter $\kappa$ where for no values of $\kappa$ the shadow of the WH falls within EHT bound. (c) The figure shows the shadow radius of Damour-Solodukin BH as a function of the parameter $\kappa$, where for all $\kappa$ values the shadow lies within the shaded region.
5.1 In vacuum
Using the above constraint, we compare the shadow results obtained for various subclasses of spacetimes in the previous section in the absence of plasma and comment on which of these can be possible candidates to represent Sgr $\rm A^{*}$. For a Schwarzschild BH, the dimensionless shadow radius is $x_{\text{sh}}=3\sqrt{3}$, which lies comfortably within the derived bounds. This aligns with the EHT analysis, where a deviation parameter of $\delta=-0.08\pm 0.09$ was reported [75]. On the other hand, the Schwarzschild wormhole, with $x_{\text{sh}}=2$, falls well outside the allowed range, disqualifying it as a viable model for Sgr A.
For the Damour-Solodukhin wormhole, the permitted range of the parameter $\sigma$ predominantly corresponds to double-barrier potentials i.e. with $\sigma$ close to 1, as illustrated in Fig.(16(a)). In case of the Hayward wormhole, the shadow radius lies significantly below the observational constraints, as shown in Fig.(16(b)). Conversely, for the regular Hayward BH, the shadow radius across the complete parameter space remains within the observational bounds, as depicted in Fig.(16(c)).
Finally, for the most general case of Hayward-Damour-Solodukhin wormhole, we identify the region in the parameter space where the shadow radius complies with the EHT constraints as shown in Fig.(17).
In summary, our analysis of the shadow radius indicates that among wormhole models, those exhibiting double- or triple-barrier effective potentials are more favored over their single-barrier counterparts. For regular BH models, current EHT observations of Sgr $\rm A^{*}$ do not impose additional constraints beyond those already provided by theoretical considerations. These conclusions may be revisited as more observational data become available in the future.
Figure 17: The allowed range of parameters of Hayward-Damour-Solodukin WH for which the shadow radius lies within the bounds obtained from Sgr $\rm A^{*}$
5.2 With plasma
We now introduce a plasma medium and examine how the bounds on both spacetime parameters and plasma parameters are affected. In the case of a Schwarzschild BH, the allowed homogeneous plasma parameter ($k_{0}$) range remains relatively small, with support only for small values of the plasma parameter as shown in Fig.(18(a)). For the non-homogeneous plasma case, the permissible parameter range is larger compared to the homogeneous scenario, however, the plasma parameter $k_{x}$ is still constrained to small values as depicted in Fig.(18(b)).
Figure 18: The figure shows the shadow radius of Schwarzschild BH in presence of (a) homogeneous and (b) non-homogeneous plasma.
Turning to the Schwarzschild wormhole, as discussed in Sec.(4.2), the shadow radius in the homogeneous plasma is independent of the plasma parameter. Consequently, the shadow remains below the observational bounds and, thus, is inconsistent with EHT results. For the non-homogeneous plasma case, Fig.(3) clearly shows that the shadow radius decreases with increasing plasma profile parameter, further deviating from the allowed range. Therefore, both plasma scenarios for the Schwarzschild wormhole are ruled out.
For the Damour-Solodukhin wormhole, the permissible range of the parameter $\sigma$ is sensitive to the plasma parameter in both homogeneous and non-homogeneous cases. In the homogeneous case, single-barrier potentials are supported for relatively large plasma parameters, whereas double-barrier potentials are favored at smaller values, as illustrated in Fig.(19(a)). In contrast, as shown in Fig.(19(b)) under a non-homogeneous plasma profile, the parameter space corresponding to single-barrier potentials is entirely excluded. However, double-barrier potentials are allowed.
Figure 19: The figure shows the allowed parameter space of Damour-Solodukin WH with (a) homogeneous plasma and (b) non-homogeneous plasma for which the shadow radius falls within the range of Sgr $\rm A^{*}$ shadow as observed by EHT .
Similar to the Schwarzschild wormhole, the Hayward wormhole’s shadow radius remains unaffected when homogeneous plasma is introduced (as discussed in Sec.(4.4)), and thus still lies below the observationally allowed range. For the non-homogeneous case, the shadow radius decreases with increasing plasma parameter, as shown in Fig.(8(a)), which further excludes this model from compatibility with EHT constraints.
We now consider the regular Hayward BH. The allowed parameter space for both homogeneous and non-homogeneous plasma is shown in Fig.(20(a)) and Fig.(20(b)), respectively. These figures indicate that the inclusion of observational constraints significantly narrows the viable range of plasma parameters initially obtained from theoretical limitations.
Figure 20: The figure shows the allowed parameter space of Hayward BH with (a) homogeneous plasma and (b) non-homogeneous plasma for which the shadow radius falls within the range of Sgr $\rm A^{*}$ shadow as observed by EHT .
Figure 21: The figure shows represent the variation of shadow radius as a function of plasma parameters for the Hayward-Damour-Solodukin wormhole in the presence of (a) homogeneous and (b) non-homogeneous plasma. The shaded regions represents the constraints on shadow radius obtained from Sgr $\rm A^{*}$.
Finally, while examining the most general case of Hayward-Damour-Solodukhin wormhole, we focused on select parameter configurations that represent distinct wormhole structures characterized by the number of peaks in their effective potential. The wormholes with values of $\sigma$ closer to 1 are characterized by multi-peak potentials as shown in Fig.(13(a)). The corresponding results for homogeneous and non-homogeneous plasma profiles are presented in Fig.(21(a)) and Fig.  (21(b)). As observed earlier, for the homogeneous plasma case, single-peak potentials are favored at higher plasma parameter values, while double- and triple-peak configurations are supported at lower values. In the non-homogeneous case, single-peak configurations are disallowed altogether, and even higher-order peak solutions are only valid within a very narrow window at small plasma parameter values.
To summarise, the Schwarzschild BH remains consistent with EHT data in both vacuum and plasma scenarios, though only for small values of the plasma parameter. In retrospect, this is expected given the small deviation parameter as mentioned in EHT results [3]. In contrast, the Schwarzschild wormhole is ruled out entirely, as its shadow radius lies below the observational lower bound and decreases further in the presence of plasma. Among wormhole spacetimes, the Damour-Solodukhin wormhole remains a viable solution. In homogeneous plasma, both single- and double-peak potential configurations can be accommodated, while in the non-homogeneous case, only double-peak configurations remain possible. A similar trend is observed for the Hayward-Damour-Solodukhin wormhole, where multi-peak potentials are favored at lower plasma parameter values, and single-peak solutions are excluded in the presence of non-homogeneous plasma. The Hayward wormhole is not supported by the EHT observations, as its shadow radius consistently falls below the allowed range regardless of plasma presence. On the other hand, the regular Hayward BH remains compatible with observations, although the inclusion of plasma significantly restricts the viable range of parameters.
6 Conclusion